A ladder leen again a perfectly smooth vertical wall at angle of 30 to the horizontal.,

A load of 800N placed 3/4 of the way up, the ladder if the ladder rest at a rough horizontal surfaces which prellent shipping. Find the magnitude and direction of the reaction between the ladder and the ground.


1
Expert's answer
2022-03-10T10:18:26-0500

The total force at the ground,


"N_g=400+800=1200(N)"


The reaction at the wall,


"\Sigma M_{gr}=0:-R\cdot l\sin30°+800\cdot\frac{1}{4}l\cos30°+400\cdot \frac{1}{2}l\cos30°=0"


"R\cdot l\sin30°=800\cdot\frac{1}{4}l\cos30°+400\cdot \frac{1}{2}l\cos30°"


"R=\frac{800\cdot\frac{1}{4}\cos30°+400\cdot \frac{1}{2}\cos30°}{\sin30°}=693(N)"